# Streaming Max-Cut in General Metrics

### v1

@@ 5 Lower Bound for Dynamic Streams
Recall that $\mathcal{X}$ is the event that the algorithm queries the distance oracle with two distinct points in $R_{i^{*}}$ during the insertion stage. The event $\mathcal{X}$ depends on $U$ and $i^{*}$, but is independent of $j^{*}$ and $K$. Recall that $\mathcal{D}$ denotes the input distribution. Our goal is to bound $\Pr_{U,i^{*}}[\mathcal{X}]$.
[…]
Fix a column $t$. Let $V\subseteq U$ be the set of IDs of $C_{t}$, and $W\subseteq U$ be the set of IDs of $P\setminus C_{t}$. Let $I$ be the internal memory state of the algorithm before the insertion of $C_{t}$, represented as an $m$-bit string. Once $V$ and $I$ are fixed, the sequence of distance oracle queries made during the insertion of $C_{t}$ is completely determined. Call this sequence $S_{1}$. Note that $S_{1}$ may be empty. In the next paragraph, we will map $S_{1}$ to $S_{3}$. We call this process refining. Analyzing $S_{3}$ suffices to bound the probability of $\mathcal{X}_{t}$.
Now we refine the query sequence. From $S_{1}$, remove all IDs belonging to $C_{t}$ to obtain $S_{2}$. Next, remove duplicates from $S_{2}$, keeping only their first occurrence. Without loss of generality, assume that $(i,j,0)$ and $(i,j,1)$ never both appear in $S_{2}$; if they do, we replace the latter occurrence with an arbitrary invalid ID. Since the IDs of $C_{t}$ are fixed, the associated invalid IDs are also predetermined, and we may freely choose one of them as the replacement. Note that in our padding scheme, such an invalid ID always takes form $(i,t,1-a_{it})$, where $1\leq i\leq n^{1/3}$ and $a_{it}=0$ or $1$. For each ID in $S_{2}$, the probability it is valid is either $1/2$ or 0, and these events are independent.
Let $S_{3}$ be the prefix of $S_{2}$ consisting of its first $2m$ IDs. If $|S_{2}|<2m$, pad $S_{3}$ with arbitrary invalid IDs. Note that $S_{1}$, $S_{2}$, and $S_{3}$ depend only on $V$ and $I$, not on $W$. Define $\mathcal{X}_{t}^{\prime}$ as the event that all IDs in $S_{3}$ are valid. Now we bound $\Pr[\mathcal{X}_{t}^{\prime}]$. If $\mathcal{X}_{t}$ occurs, then either $\mathcal{X}_{t}^{\prime}$ occurs, or $S_{3}$ contains an ID of $R_{i^{*}}$. Condition on fixed $V$ and $I$. Since each of the $2m$ IDs in $S_{3}$ is valid independently with probability at most $1/2$, we have $\Pr_{W}[\mathcal{X}_{t}^{\prime}]\leq\frac{1}{2^{2m}}$. Since $I$ has at most $2^{m}$ possible states, a union bound gives
$\Pr_{W}[\exists I:\mathcal{X}_{t}^{\prime}]\leq\frac{1}{2^{m}}.$ (14)
Suppose $\mathcal{X}_{t}$ occurs. Then, before the algorithm queries any invalid ID, it must query an ID corresponding to a point in $R_{i^{*}}\setminus\{p_{i^{*},t}\}$. This can occur in two ways: 1. $S_{3}$ contains an ID of a point in $R_{i^{*}}$, or 2. all IDs in $S_{3}$ are valid.
The first event occurs with probability at most $2m/n^{2/3}$, since $S_{3}$ has size $2m$ and $i^{*}$ is uniformly distributed in $n^{2/3}]$. The second event occurs with probability at most $1/2^{m}$, by (14). Thus, $\Pr_{i^{*},W}[\mathcal{X}_{t}]\leq\frac{1}{2^{m}}+\frac{2m}{n^{2/3}}$.
Since the above holds for every $t\in[n^{1/3}]$, we conclude
$\Pr_{i^{*},U}[\mathcal{X}]=\Pr_{i^{*},U}\!\left[\bigcup_{t=1}^{n^{1/3}}\mathcal{X}_{t}\right]\leq\sum_{t=1}^{n^{1/3}}\Pr_{i^{*},U}[\mathcal{X}_{t}]\leq n^{1/3}\left(\frac{1}{2^{m}}+\frac{2m}{n^{2/3}}\right)\leq 0.05.$
@@
Recall that $\mathcal{Y}$ is the event that the algorithm queries the distance between $p_{i^{*},j^{*}}$ and another point in $R_{i^{*}}$ during the deletion stage, $\cal D$ is the input distribution, and $x$ is the input instance. Our goal is to bound $\Pr_{T\sim\mathcal{D}}[\mathcal{Y}]$.
Let $i^{*}$ be fixed. Let $G\subseteq U$ be the set of IDs of $R_{i^{*}}$, and $H\subseteq U$ be the set of IDs of $P\setminus R_{i^{*}}$. Let $I$ be the internal memory state of the algorithm before the deletion, represented as an $m$-bit string. Once $G$ and $I$ are fixed, the sequence of distance oracle queries made during the deletion is completely determined. Call this sequence $S_{1}^{\prime}$.
Now we refine the query sequence. From $S_{1}^{\prime}$, remove all IDs belonging to $P\setminus R_{i^{*}}$ to obtain $S_{2}^{\prime}$. Next, remove duplicates from $S_{2}^{\prime}$, keeping only their first occurrence. Without loss of generality, assume that $(i,j,0)$ and $(i,j,1)$ never both appear in $S_{2}^{\prime}$; if they do, we replace the latter occurrence with an arbitrary invalid ID. For each ID in $S_{2}^{\prime}$, the probability it is valid is either $1/2$ or 0, and these events are independent.
Let $S_{3}^{\prime}$ be the prefix of $S_{2}^{\prime}$ consisting of its first $2m$ IDs. If $|S_{2}^{\prime}|<2m$, pad $S_{3}^{\prime}$ with arbitrary invalid IDs. Note that $S_{1}^{\prime}$, $S_{2}^{\prime}$, and $S_{3}^{\prime}$ depend only on $G$ and $I$, not on $H$. Define $\cal W$ as the event that all IDs in $S_{3}^{\prime}$ are valid.
[…]
$\Pr_{H}[\exists I:\mathcal{W}]\leq\frac{1}{2^{m}}.$ (15)
Suppose $\mathcal{Y}$ occurs. Then, before the algorithm queries any invalid ID, it must query the ID of $p_{i^{*},j^{*}}$. This can occur in two ways: a) $S_{3}^{\prime}$ contains the ID of $p_{i^{*},j^{*}}$, or b) all IDs in $S_{3}^{\prime}$ are valid.
The first event occurs with probability at most $2m/n^{1/3}$, since $R_{i^{*}}$ has size $n^{1/3}$ and $j^{*}$ is uniformly distributed in $n^{1/3}]$. The second event occurs with probability at most $1/2^{m}$, by (15). Thus, $\Pr_{j^{*},H}[\mathcal{Y}]\leq\frac{1}{2^{m}}+\frac{2m}{n^{1/3}}$.
Since the above holds for every $i^{*}\in[n^{2/3}]$ and $G$, we conclude
$\Pr_{i^{*},j^{*},U}[\mathcal{Y}]\leq\frac{1}{2^{m}}+\frac{2m}{n^{2/3}}\leq 0.05.$

### v2

@@ 5 Lower Bound for Dynamic Streams
During the insertion stage, the columns $C_{1},\ldots,C_{n^{1/3}}$ arrive sequentially. For each $t\in[n^{1/3}]$, let $\mathcal{X}_{t}$ denote the event that $\mathcal{X}$ occurs during the insertion of $C_{t}$, i.e., from the time the algorithm reads $\text{ID}_{t,1}$ until it is ready to read $\text{ID}_{t+1,1}$ (for $t=n^{1/3}$, this stage ends when the first deletion begins). By the union bound, $\Pr[\mathcal{X}]\leq\sum_{t=1}^{n^{1/3}}\Pr[\mathcal{X}_{t}]$. It therefore suffices to bound $\Pr[\mathcal{X}_{t}]$ for each $t$.
Fix a column $t$. Let $V\subseteq U$ be the set of IDs of $C_{t}$. Let $I$ be the internal memory state of the algorithm immediately before the insertion of $C_{t}$, represented as an $m$-bit string. Denote this query sequence of point pairs as $S$, and let $T:=\{s:(s,t)\in S\}\cup\{t:(s,t)\in S\}$ be the set of distinct point IDs that belong to any query pair (so $T$ is not a multiset). Furthermore, exclude the IDs in $C_{t}$ from $T$, i.e., let $T^{\prime}:=T\setminus V$.
Fact 5.7.
$T^{\prime}$ is determined by $V$ and $I$, and is independent of $i^{*}$.
Proof.
The fact that $T^{\prime}$ is determined by $I$ and $V$ is by definition. The randomness of $V$ comes from $U$ which is the IDs. On the other hand, since the algorithm is deterministic, $I$ depends on $\operatorname{dist}^{\prime}$ (which is deterministic) as well as the set of IDs ever inserted (whose randomness comes from $U$, and more specifically from the IDs of $C_{1},\ldots,C_{t-1}$). ∎
Lemma 5.8.
With respect to the randomness of IDs, $\Pr[|T^{\prime}|>2m]\leq\frac{1}{2^{m}}$.
Proof.
The fact that $T^{\prime}$ depends on $I$ which depends on $U$ (the IDs) makes the analysis of randomness tricky, and here we need to do a low-level counting analysis with respect to the behavior of the deterministic algorithm.
This lemma is trivial when $t=1$, since $|T^{\prime}|$ can only be 0 as we remove the IDs of $C_{1}$ which is the only input points seen so far. Next, we focus on $t\geq 2$.
We analyze the conditional probability $\Pr[|T^{\prime}|>2m\mid V]$, for any $V$, and we fix some $V$ in the following analysis. For $Z\in\{0,1\}^{m}$, define $T^{\prime}_{Z}$ be the realization of $T^{\prime}$ such that $I=Z$ (and given $V$). Since the algorithm is deterministic, we can interpret the behavior of the algorithm as two mappings: a) a mapping $f$ that maps the input IDs on $C_{1},\ldots,C_{t-1}$, denoted as $U_{t-1}$, to the internal state $I$, i.e., $f:U_{t-1}\mapsto I$; and b) a mapping $g$ that maps the internal state $I$ to $T^{\prime}_{I}$, i.e., $g:I\mapsto T^{\prime}_{I}$. Hence,
$\Pr[|T^{\prime}|>2m\mid V]=\sum_{Z\in\{0,1\}^{m}}\mathbb{I}[|T^{\prime}_{Z}|>2m]\cdot\frac{|f^{-1}(Z)|}{2^{|U_{t-1}|}},$ (20)
where $\mathbb{I}$ is the indicator function. If $|T^{\prime}_{Z}|>2m$, we take the $2m$-length prefix of $T^{\prime}_{Z}$, denoted as $T^{\prime\prime}_{Z}$, and this $T^{\prime\prime}_{Z}\subseteq U_{t-1}$ defines a specific realization of IDs $(i,j,a_{i,j})$ of some $2m$ positions $(i,j)$’s in $C_{1},\ldots,C_{t-1}$. Note that for different $Z$ this $T^{\prime\prime}_{Z}$ may also be different.
Now, denote the subset of $f^{-1}(Z)$ that satisfies a realization $T^{\prime\prime}_{Z}$ as $F_{Z,T^{\prime\prime}_{Z}}$. Then $\frac{|F_{Z,T^{\prime\prime}_{Z}}|}{2^{|U_{t-1}|}}$ is the probability that we see this specific realization. Therefore, continue from (20)
$\Pr[|T^{\prime}|>2m\mid V]=\sum_{Z\in\{0,1\}^{m}}\mathbb{I}[|T^{\prime}_{Z}|>2m]\cdot\frac{|f^{-1}(Z)|}{2^{|U_{t-1}|}}\leq\sum_{Z\in\{0,1\}^{m}}\frac{|F_{Z,T^{\prime\prime}_{Z}}|}{2^{|U_{t-1}|}}.$ (21)
Observe that $F_{Z,T^{\prime\prime}_{Z}}$ is a subset of the $2^{|U_{t-1}|}$ realizations of IDs in $U_{t-1}$, such that a fixed $2m$ locations of $(i,j)\in C_{1}\cup\ldots\cup C_{t-1}$’s takes a fixed realization in $\{0,1\}^{2m}$, as specified in $T^{\prime\prime}_{Z}$. Therefore, at most $|U_{t-1}|-2m$ locations are free from the constraint and $|F_{Z,T^{\prime\prime}_{Z}}|\leq 2^{|U_{t-1}|-2m}$. Substituting this to (21), we have
$\Pr[|T^{\prime}|>2m\mid V]\leq\sum_{Z\in\{0,1\}^{m}}\frac{|F_{Z,T^{\prime\prime}_{Z}}|}{2^{|U_{t-1}|}}\leq\sum_{Z\in\{0,1\}^{m}}\frac{1}{2^{2m}}=\frac{1}{2^{m}}.$
We finish the proof of Lemma 5.8 by taking $\Pr[|T^{\prime}|>2m]=\mathbb{E}[\Pr[|T^{\prime}|>2m\mid V]]\leq\frac{1}{2^{m}}$. ∎
Now, condition on the event that $|T^{\prime}|\leq 2m$, since $T^{\prime}$ is independent of $i^{*}$ (Fact 5.7), we conclude that $\Pr[\mathcal{X}_{t}\mid|T^{\prime}|\leq 2m]\leq\frac{2m}{n^{2/3}}$, where we are using that $i^{*}$ is chosen uniformly at random from $n^{2/3}]$ and $T^{\prime}$ has $2m$ places to hit $i^{*}$. Hence, by Lemma 5.8
$\Pr[\mathcal{X}_{t}]\leq\Pr[\mathcal{X}_{t}\mid|T^{\prime}|\leq 2m]+\Pr[|T^{\prime}|>2m]\leq\frac{1}{2^{m}}+\frac{2m}{n^{2/3}}.$
This further implies that
$\Pr[\mathcal{X}]\leq\sum_{t=1}^{n^{1/3}}\Pr[\mathcal{X}_{t}]\leq n^{1/3}\left(\frac{1}{2^{m}}+\frac{2m}{n^{2/3}}\right)\leq 0.05.$ (22)
@@
Recall that $\mathcal{Y}$ is the event that the algorithm queries $\operatorname{dist}^{\prime}$ between $p_{i^{*},j^{*}}$ and another point in $R_{i^{*}}$ during the deletion stage. The proof is similar to that for insertion-only, and we only sketch the key steps and the differences.
Condition on $i^{*}$, and we use the randomness of $j^{*}$ and the IDs, i.e., we analyze $\Pr[\mathcal{Y}\mid i^{*}]$. Let $G\subseteq U$ be the set of IDs of $P\setminus R_{i^{*}}$, and let $I$ be the internal memory state of the algorithm before the deletion, represented as an $m$-bit string.
Similar to the insertion stage, for each realization $Z\in\{0,1\}^{m}$ of the internal states of the algorithm, we define $T$ as the distinct point IDs that belong to any query pair, and let $T^{\prime}:=T\setminus G$ be that excluding IDs in $G$. We can establish a similar lemma as in Lemma 5.8: we have $\Pr[|T^{\prime}|>2m\mid i^{*}]\leq\frac{1}{2^{m}}$, which follows from the randomness of the IDs.
Next, observe that $T^{\prime}$ is independent of $j^{*}$, and we have $\Pr[\mathcal{Y}\mid|T^{\prime}|\leq 2m,i^{*}]\leq\frac{2m}{n^{1/3}}$, where we are using the randomness of $j^{*}$, which is chosen uniformly at random from $n^{1/3}]$ and $T^{\prime}$ has $2m$ places to hit $j^{*}$. Hence,
$\Pr[\mathcal{Y}\mid i^{*}]\leq\Pr[\mathcal{Y}\mid|T^{\prime}|\leq 2m,i^{*}]+\Pr[|T^{\prime}|>2m\mid i^{*}]\leq\frac{1}{2^{m}}+\frac{2m}{n^{2/3}}\leq 0.05.$
Therefore, $\Pr[\mathcal{Y}]\leq\mathbb{E}[\Pr[\mathcal{Y}\mid i^{*}]]\leq 0.05$.